Hi. I'm trying to figure out how many employees a company needs to add base on the sales volume increased? 40 hours work week.

Sales Revenue=$1,020,000
Current Employees=74
Total hours for the month=13,757
Overtime dollar percentage=26%
Overtime dollar=$63,000

Would you be able to assist me in finding out how many more workers needed based on the above info? I don't know what formula to use?

I was thinking $1,020,000/74=$13,783 per ee / 26% = $53,014 / 160 = $331 / 8 = 41 employee?

Thanks,
Wen
1
Expert's answer
2016-06-10T09:59:09-0400
40 hours work week, Sales Revenue = $1,020,000, Current Employees = 74, Total hours for the month = 13,757, Overtime dollar percentage = 26%, Overtime dollar = $63,000.
If a worker should work 40 hours a week, then he should work 40*4 = 160 hours a months.
If total hours for the month are 13,757, then we need 13,757/160 = 85.98 = 86 workers to work without the overtime.
So, we need 86 - 74 = 12 more workers to hire.